# The Automorphism Equivariant Hitchin Index

### v1

@@ 3 Localization
which easily follows from the fact that $\Phi_{p}(\hat{\zeta}^{u})=0$ in $S^{-1}\cdot\mathbb{Z}[\mu_{p}]$ for $u=1,...,p-1$. Multiplying by $p$ and using equation (94) we obtain the following identity in $K_{\mathbb{T}\times\mu_{p}}^{0}(\mathcal{M}_{i,j})/\ker(L)$
@@
$p\cdot T_{i,j}\mathcal{M}=p\cdot T_{i,j}\mathcal{M}_{i}\cdot\Phi_{p}(\hat{\zeta})+V_{i,j}-\pi_{0}(V_{i,j})\cdot\Phi_{p}(\hat{\zeta}).$ (96)
We can now apply $\omega$ to equation (96). Using that $\omega$ is a homomorphism that transforms sums into products, we obtain
@@
Let $u$ be a formal variable. Denote by $s_{u}=\sum_{m\geq 0}u^{m}s^{m}$ the operation of taking the graded sum of symmetric powers, and denote by $\lambda_{u}=\sum_{m\geq 0}u^{m}\lambda^{m}$ the operation of taking the graded sum of exterior powers. Recall that for any equivariant vector bundle $V$ we have that $s_{u}(V)\lambda_{-u}(V)=1$. Therefore we introduce the notation $\lambda(W)^{-1}=s(W)$ for any equivariant bundle $W$ with trivial $\mathbb{T}$-action for which this class exists in $S^{-1}\cdot K^{0}_{\mathbb{T}\times\langle f\rangle}(\mathcal{M}_{i,j})$. By Nielsen’s localization theorem (Theorem 3.3), this holds for $N^{*}_{i,j}$. We have
$s\left(N^{*}_{i,j}\right)=\lambda(T^{*}\mathcal{M}_{i,j})\cdot s(T^{*}_{i,j}\mathcal{M}^{\mathbb{T}}).$ (98)
Now we will rewrite the factor $s(T^{*}_{i,j}\mathcal{M}^{\mathbb{T}}).$ Using that $\pi_{0}(T^{*}_{i,j}\mathcal{M}^{\mathbb{T}})=T^{*}\mathcal{M}_{i,j}$ and the homomorphism property of $s$, we obtain from (96) the following equation
$\begin{split}s(T^{*}_{i,j}\mathcal{M}^{\mathbb{T}})^{p}&=s(T^{*}\mathcal{M}_{i,j}\cdot\Phi_{p}(\hat{\zeta}))^{p}\prod_{x\in X^{f}}\frac{s((V^{\mathbb{T}}_{x,i,j})^{*})}{s(\pi_{0}(V^{\mathbb{T}}_{x,i,j})^{*}\cdot\Phi_{p}(\hat{\zeta}))}.\end{split}$ (99)
We now observe that for all equivariant vector bundles $W\rightarrow\mathcal{M}_{i,j}$, for which $s(W)$ exists, we have that
$\Omega(W)=\omega(W)\cdot s((W^{\mathbb{T}}_{i,j})^{*}).$ (100)
By multiplying by $\lambda(N^{*}_{i,j})^{-p}$ on both sides of equation (97), and using equation (98), equation (99), equation (100) and the definition of $\tilde{\chi}_{i,j}$ given in equation (8), we obtain the following identity
$\omega(T_{i,j}\mathcal{M})^{p}\cdot\lambda\left(N^{*}_{i,j}\right)^{-p}=\varXi_{i}^{*}(\tilde{\chi}_{i,j})^{p}.$ (101)
Since $p$ is odd, it is trivial to see that equation (96) implies that $(-1)^{\nu_{i,j}}=(-1)^{\tilde{\nu}_{i,j}}.$ Thus it only remains to extract the correct $p$’th root. For each $l=1,...,p-1$ define $a_{i,j,l}=\operatorname{rank}(\pi_{l}(N^{*}_{i,j}))$, and define $b_{i,j,l}=\operatorname{rank}(\pi_{l}(T_{i,j}\mathcal{M}^{-}))$. Further, define $\rho_{i,j}=\sum_{l=1}^{p-1}b_{i,j,l}\cdot l.$ Consider the projection, which takes $\operatorname{ch}_{0}$ of the leading term of a $1/t$-series
@@
Upon inspecting the definition of $\omega$ one finds that
$P_{0}\left(\operatorname{ch}(\omega(T_{i,j}\mathcal{M})\cdot\lambda\left(N^{*}_{i,j})\right)^{-1}\right)=\hat{\zeta}^{\rho_{i,j}}\prod_{l=1}^{p-1}(1-\hat{\zeta}^{l})^{-a_{i,j,l}}.$ (103)
@@
$p\left(T_{i,j}\mathcal{M}^{\mathbb{T}}-T\mathcal{M}_{i,j}\right)=p\sum_{l=1}^{p-1}T\mathcal{M}_{i,j}\cdot\hat{\zeta}^{l}+V^{\mathbb{T}}_{i,j}-\pi_{0}(V_{i,j}^{\mathbb{T}})\cdot\Phi_{p}(\hat{\zeta}),$ (104)
$pT_{i,j}\mathcal{M}^{-}=p(T_{i,j}\mathcal{M}_{i})^{-}\cdot\Phi_{p}(\hat{\zeta})+V_{i,j}^{-}-\pi_{0}(V_{i,j})^{-}\cdot\Phi_{p}(\hat{\zeta}),$ (105)
@@
Thus it is clear that $a_{i,j,l}=\tilde{a}_{i,j,l}$ and $b_{i,j,l}=\tilde{b}_{i,j,l}$. Therefore also $\rho_{i,j}=\tilde{\rho}_{i,j}$. It follows that the right hand side of equation (103) is equal to $\tilde{\zeta}_{i,j}$. Thus we can conclude by applying Lemma 3.2 applied to the elements $a=\operatorname{ch}(\tilde{\chi}_{i,j})$ and $b=\operatorname{ch}(\omega(T_{i,j}\mathcal{M})\cdot\lambda(N^{*}_{i,j})^{-1}).$
@@ 3.1 The Galois Case
Recall the notation $\pi_{\mathcal{M}}:\mathcal{M}\times X\rightarrow\mathcal{M}$ for the natural projection onto the first factor. Assume that $X^{f}=\emptyset$. Then Hurwitz formula specialises to $\chi(X)=p\chi(\tilde{X})$, where $\chi(X)$ (resp. $\chi(\tilde{X}))$) denotes the Euler characteristic of $X$ (resp. $\tilde{X}$). Fix $\tilde{x}_{0}\in\tilde{X}$ and define
@@
Proof of Corollary 1.2.
In the case $X^{f}$ is empty, pullback with respect to the projection $\pi:X\rightarrow\tilde{X}$ induces an isomorphism of $\mathbb{T}$-varities $\pi^{*}:\tilde{\mathcal{M}}\rightarrow\mathcal{M}^{f}$. We write $\varXi=(\pi^{*})^{-1}$ for the inverse isomorphism of $\mathbb{T}$-varieties. We have that $(\varXi^{-1})^{*}(\mathcal{L})\cong\tilde{\mathcal{L}}^{\otimes p}$. We simplify the notation and write $\{\tilde{F}_{a}\}_{a\in A}$ for the components of ~$\tilde{\mathcal{M}}^{\mathbb{T}}$. Let $a\in A$. Denote by $T_{a}\tilde{\mathcal{M}}$ the restriction of $T\tilde{\mathcal{M}}$ to $\tilde{F}_{a}$ and write ${\nu}_{a}=\operatorname{rank}(T_{a}\tilde{\mathcal{M}}^{-})$. Observe that as $\operatorname{\tilde{\mathcal{D}}}$ is empty, we have that $\tilde{\nu}_{a}=\nu_{a}$. Define the $t^{-1}$-series
@@
Then $\chi_{t}(\tilde{\mathcal{M}},\tilde{\mathcal{L}}^{k})(t)=\sum_{a\in A}\chi_{a}(t)$ by the localization result of Wu [68, Theorem 3.14] recalled in Theorem 3.1 in this article. By Theorem 1.2 it will therefore be sufficient to prove that for all $a\in A$ we have that
@@
We now analyze the left hand side of equation (110). Recall the notation $s=\lambda^{-1}$. As $\operatorname{\tilde{\mathcal{D}}}$ is empty, we find that
$\tilde{\chi}_{a}=\omega(T_{a}\tilde{\mathcal{M}}\cdot\Phi_{p}(\hat{\zeta}))\prod_{j=1}^{p-1}s(T^{*}\tilde{F}_{a}\cdot\hat{\zeta}^{j}).$ (111)
@@
$\operatorname{Td}\tilde{F}_{a}=\exp(c_{a}/2)\prod_{z}\frac{2}{e^{z/2}-e^{-z/2}}.$ (113)
@@
$\prod_{j=1}^{p-1}\operatorname{ch}(s(T^{*}\tilde{F}_{a}\cdot\hat{\zeta}^{j}))=\prod_{\alpha\in\mu_{p}^{*}}\prod_{z}\frac{1}{1-\alpha e^{-z}}=\frac{\exp\left((p-1)c_{l}/2\right)}{\prod_{\alpha\in\mu_{p}^{*}}\prod_{z}\left(e^{z/2}-\alpha e^{-z/2}\right)}$ (114)
@@
$Z^{p}-Z^{-p}=\prod_{\alpha\in\mu_{p}}(Z-\alpha Z),$ (116)
@@
$\prod_{j=1}^{p-1}\operatorname{ch}(s(T^{*}\tilde{F}_{a}\cdot\hat{\zeta}^{j}))\cdot\operatorname{Td}\tilde{F}_{a}=p^{-d_{l}}\operatorname{Td}\tilde{F}_{a}(pz).$ (117)
@@ 4 Quantum Topology
Proof of Corollary 1.3.
@@
It is thus clear that $(D_{\Lambda},l_{0})$ is determined by the Seifert invariants of $(M_{f},K)$ as defined in Section 1.3. ∎
We now explain the refinement formula given in equation (17). We have that $\mathcal{N}^{f}$ is a subvariety of $\mathcal{M}^{\mathbb{T}\times\langle f\rangle}$. The torus $\mathbb{T}$ act trivially on the restriction of the determinant line bundle $\mathcal{L}$ to $\mathcal{N}$, whereas $\mathbb{T}$ acts with non-trivial weight $t^{(2-2g)m}$ on $\mathcal{L}$ restricted to $\mathcal{F}_{m}$ for each $m\in M$. Therefore the sum appearing in (9) corresponding to components of $\mathcal{N}^{f}$ is seen to give the constant term of $\chi_{\mathbb{T}}(\mathcal{M},\mathcal{L}^{k},f)$. We have that $T_{\mathcal{N}}\mathcal{M}^{-}\cong T\mathcal{N}$, and $\det(T\mathcal{N})\cong\mathcal{L}^{2}$. As the rank of $T_{\mathcal{N}}M^{-}$ is equal to $3g-g$, the constant is equal to $(-1)^{3g-3}$ times $\chi(\mathcal{N},\mathcal{L}^{k+2},f)$. Thus we obtain equation (17).

### v2

@@ 2.1 Application of Localization
which easily follows from the fact that $\Phi_{p}(\hat{\zeta}^{u})=0$ in $S^{-1}\cdot\mathbb{Z}[\mu_{p}]$ for u=1,,p1u=1,...,p-1. Multiplying by $p$ and using equation (106) we obtain the following identity in $K_{\mathbb{T}\times\mu_{p}}^{0}(\mathcal{M}_{i,j})/\ker(L)$
@@
[…]
We can now apply $\Omega$ to equation (108). Using that $\Omega$ is a homomorphism that transforms sums into products, we obtain the desired consequence of equation (100)
@@
Since $p$ is odd, it is trivial to see that equation (108) implies that $(-1)^{\nu_{i,j}}=(-1)^{\tilde{\nu}_{i,j}}.$ Thus it only remains to extract the correct $p$’th root. Let $N^{*}_{i,j}$ be the conormal of $\mathcal{M}_{i,j}\subset\mathcal{M}^{\mathbb{T}}$. For each l=1,,p1l=1,...,p-1 define $a_{i,j,l}=\operatorname{rank}(\pi_{l}(N^{*}_{i,j}))$, and define $b_{i,j,l}=\operatorname{rank}(\pi_{l}(T_{i,j}\mathcal{M}^{-}))$. Further, define $\rho_{i,j}=\sum_{l=1}^{p-1}b_{i,j,l}\cdot l.$ Consider the projection, which takes $\operatorname{ch}_{0}$ of the leading term of a $1/t$-series
@@
Upon inspecting the definition of $\omega$ and using (77) one finds that
$\begin{split}P_{0}(\operatorname{ch}(\Omega(T_{i,j}\mathcal{M})))&=P_{0}\left(\operatorname{ch}(\omega(T_{i,j}\mathcal{M})\cdot\lambda\left(N^{*}_{i,j})\right)^{-1}\right)\\ &=\hat{\zeta}^{\rho_{i,j}}\prod_{l=1}^{p-1}(1-\hat{\zeta}^{l})^{-a_{i,j,l}}.\end{split}$ (111)
@@
[…]
@@
Thus it is clear that $a_{i,j,l}=\tilde{a}_{i,j,l}$ and $b_{i,j,l}=\tilde{b}_{i,j,l}$. Therefore also $\rho_{i,j}=\tilde{\rho}_{i,j}$. It follows that the right hand side of equation (111) is equal to $\tilde{\zeta}_{i,j}$. Thus we can conclude by applying Lemma 2.2 applied to the elements $a=\operatorname{ch}(\tilde{\chi}_{i,j})$ and $b=\operatorname{ch}(\Omega(T_{i,j}\mathcal{M})).$
Second part.
We now prove (101). Define a $\mathbb{T}$-equivariant line bundle on $\tilde{\mathcal{M}}_{i}$ by $\check{\mathcal{L}}_{i}=(\varXi_{i}^{-1})^{*}(\mathcal{L}_{\mid\mathcal{M}_{i}})$. We must show that $c_{1}(\check{\mathcal{L}}_{i})=\tilde{\omega}_{i}$. Recall that $\tilde{\mathcal{N}}_{i}$ is diffeomorphic to a certain moduli space of flat $G$-bundles on $\tilde{X}\setminus\operatorname{\tilde{\mathcal{D}}}_{i}$. Through this identification, the moduli space $\tilde{\mathcal{N}}_{i}$ supports an induced symplectic form $\tilde{\Omega}_{i}$ as constructed in [32], whose cohomology class satisfy
$\tilde{\Omega}_{i}]=\left(\tilde{\alpha}_{i}+\sum_{x\in\operatorname{\tilde{\mathcal{D}}}_{i}}w_{i,2}(x)\delta_{i,x}\right)_{\mid\mathcal{N}_{i}}=\frac{\tilde{\omega_{i}}_{\mid\tilde{\mathcal{N}}_{i}}}{p}.$ (116)
It was observed in [3] that
$c_{1}(\check{\mathcal{L}}_{i\mid\tilde{\mathcal{N}}_{i}})=\tilde{\omega}_{i\mid\tilde{\mathcal{N}}_{i}}=p[\tilde{\Omega}_{i}].$ (117)
The natural map $\operatorname{Pic}(\tilde{\mathcal{M}}_{i})\rightarrow\operatorname{Pic}(\tilde{\mathcal{N}}_{i})$ is an isomorphism. This is proven in [67] in the case of moduli spaces of stable parabolic Higgs bundles of coprime rank and degree and stable parabolic holomorphic vector bundles of coprime rank and degree, and the result generalizes easily to the cased of fixed determinant. Therefore, equation (117) implies the following identity in cohomology
$c_{1}(\check{\mathcal{L}}_{i})=\tilde{\omega}_{i}.$ (118)
Thus it only remains to argue that this identity holds as an identity in $\mathbb{T}$-equivariant cohomology.
Recall that $\mathbb{T}$-equivariant cohomology of a point is additively isomorphic to a polynomial ring of one formal variable, which we denote by $u$, of cohomological degree two. Recall that as $\mathcal{M}$ is a semi-projective variety in the sense of [23], we have that $H^{*}(\mathcal{M})\cong H^{*}(\mathcal{M}^{\mathbb{T}})$ and by [53] the $\mathbb{T}$-equivariant cohomology ring $H^{*}_{\mathbb{T}}(\mathcal{M})$ is isomorphic as a $\mathbbm{Q}[u]$-module to the ring of polynomials in $u$ with coefficients in $H^{*}(\mathcal{M})$, i.e. in symbols
$H^{*}_{\mathbb{T}}(\mathcal{M})\cong H^{*}(\mathcal{M})[u]\cong H^{*}(\tilde{\mathcal{M}}^{\mathbb{T}})[u].$ (119)
Thus, for any $\mathbb{T}$-equivariant line bundle $L\rightarrow\mathcal{M}$, and any component $Z\subset\mathcal{M}^{\mathbb{T}},$ the weight of the $\mathbb{T}$-action on $L_{\mid Z}$ is equal to an integer $b$ if and only if the coefficient of $u$ in $c_{1}(\mathcal{L})_{\mid Z}$ is equal to $b$. Similarly, $H^{*}_{\mathbb{T}}(\tilde{\mathcal{M}}_{i})\cong H^{*}(\tilde{\mathcal{M}}_{i}^{\mathbb{T}})[u]$. Because of this isomorphism, we see that establishing (101) for every $i\in\mathcal{I}$ and $j\in\mathcal{J}_{i}$ is equivalent to establing (LABEL:eq:identificationofinducedTLibebundle) for every $i\in\mathcal{I}$.
Thus, we mush show that for every component $\tilde{\mathcal{M}}_{i,j}$ of $\tilde{\mathcal{M}}_{i}^{\mathbb{T}}$ the following holds in $\mathbb{T}$-equivariant cohomology
$c_{1}(\check{\mathcal{L}}_{i\mid\tilde{\mathcal{M}}_{i,j}})=\tilde{\omega}_{i\mid\tilde{\mathcal{M}}_{i,j}}.$ (120)
Because of the identity between complex line bundles given in equation (118), we see that equation (120) is equivalent to the coefficient of $u$ being the same in $c_{1}(\mathcal{L})_{\mid\mathcal{M}_{i,j}}$ and in $\tilde{\omega}_{i\mid\tilde{\mathcal{M}}_{i,j}}$. This fact was already established for the component corresponding to $\tilde{\mathcal{N}}_{i}$ (as the coefficient of $u$ is equal to zero on both sides of the equation). Thus, it remains to prove this fact for a component $\mathcal{M}_{i,j}$ with $j\in\mathcal{J}_{i,c}$ and $c\in C$.
Write $\alpha=c_{1}(\mathcal{L}).$ By definition, we have that $\mathcal{M}_{i,j}\subset\mathcal{F}_{c}$, and by [45, Lemma 6.1] it holds that the coefficient of $\alpha_{\mid\mathcal{F}_{c}}$ is equal to $\chi+c$, where $\chi=2-2g$ is the Euler characteristic of the Riemann surface $X$. By adapting the proof of [45, Lemma 6.1] one can show that the coefficient of $u$ in the restriction of $\tilde{\alpha}_{i}$ to $\tilde{\mathcal{M}}_{i,j}$ is equal to the integer $\mu_{ij}$ defined to be
$\mu_{ij}=\tilde{\chi}-\deg(\mathcal{\operatorname{\tilde{\mathcal{D}}}}_{i})+\deg(\operatorname{\tilde{\mathcal{D}}}_{ij})+l_{j},$ (121)
where $\tilde{\chi}=2-2\tilde{g}$ is the Euler characteristic of $\tilde{X}$, $l=l_{j}=(c-\deg(j))/p$ and $\operatorname{\tilde{\mathcal{D}}}_{ij}$ is the divisor defined in (71). Similarly, the coefficient of $u$ in $\delta_{i,x}$ restricted to $\tilde{\mathcal{M}}_{i,j}$ is equal to $\epsilon_{j,x}$. Recall $pw_{i,2}(x)=b_{i}(x)$ for each $x\in\operatorname{\tilde{\mathcal{D}}}_{i}$ by the definitions of $w_{i}$ and $b_{i}$. Hence, the coefficient $c_{ij}$ of $u$ in $\tilde{\omega}_{i}$ restricted to $\tilde{\mathcal{M}}_{i,j}$ is equal to
$c_{i,j}=\mu_{i,j}+\sum_{x\in\operatorname{\tilde{\mathcal{D}}}_{i}}\epsilon_{j,x}b_{i}(x),$ (122)
and we must show that $c_{i,j}=\chi+c.$
Recall that since $X\rightarrow\tilde{X}$ is a branched covering of prime order $p$, Hurwitz’ formula simplifies to
$\chi=p\tilde{\chi}-(p-1)\lvert X^{f}\rvert$. (123)
Using these facts, the expression for $p\deg(\operatorname{\tilde{\mathcal{D}}}_{ij})=\deg(\pi^{*}(\operatorname{\tilde{\mathcal{D}}}_{ij}))$ given by the right hand side of equation (71), and the fact that $pl_{j}+\deg(j)=c$ by definition of $j\in\mathcal{J}_{i,c}$, we see that the coefficient $c_{ij}$ of $u$ in $\tilde{\omega}_{i}$ restricted to $\tilde{\mathcal{M}}_{i,j}$ is equal to $\chi+c$ by the following computation
$c_{ij}$ $=p\mu_{ij}+\sum_{x\in\operatorname{\tilde{\mathcal{D}}}_{i}}b_{i}(x)\epsilon_{j,x}$ (124)
$=p(\tilde{\chi}-\deg(\mathcal{\operatorname{\tilde{\mathcal{D}}}}_{i})+\deg(\operatorname{\tilde{\mathcal{D}}}_{ij})+l)+\sum_{x\in\operatorname{\tilde{\mathcal{D}}}_{i}}b_{i}(x)\epsilon_{j,x}$ (125)
[…]
$+\sum_{x\in\operatorname{\tilde{\mathcal{D}}}_{i}}b_{i}(x)\epsilon_{j,x}=(p\tilde{\chi}-(p-1)\lvert X^{f}\rvert)+(pl+\deg(j))=\chi+c.$ (127)
This finishes the proof. ∎
@@ 2.2 The Galois Case
[…]
@@
With notation as in the introduction, let $\tilde{\mathcal{M}}$ be the moduli space of stable Higgs bundles on $\tilde{X}$ of rank two and fixed determinant $\tilde{\Lambda}$ of odd degree. Let $\tilde{\operatorname{\mathbbm{E}}}\rightarrow\tilde{\mathcal{M}}\times\tilde{X}$ be the universal Higgs bundle. Consider the projection $\pi_{\tilde{\mathcal{M}}}:\tilde{\mathcal{M}}\times\tilde{X}\rightarrow\tilde{\mathcal{M}}$ and define
$\tilde{\mathcal{L}}=\det(\tilde{\operatorname{\mathbbm{E}}}_{\mid\tilde{\mathcal{M}}\times\{x_{0}\}})^{1-\tilde{g}}\otimes\det((\pi_{\tilde{\mathcal{M}}})_{!}(\tilde{\operatorname{\mathbbm{E}}})[1]),$ (129)
where $(\pi_{\tilde{\mathcal{M}}})_{!}(\tilde{\operatorname{\mathbbm{E}}})[1]$ is the $1$-shift of the derived pushforward of $\tilde{\operatorname{\mathbbm{E}}}$.
Proof of Corollary LABEL:thm:Galois.
[…]
@@
Then $\chi_{t}(\tilde{\mathcal{M}},\tilde{\mathcal{L}}^{k})(t)=\sum_{a\in A}\chi_{a}(t)$ by the localization result of Wu [79, Theorem 3.14] recalled in Theorem 2.1 in this article. By Theorem LABEL:ThmMain it will therefore be sufficient to prove that for all $a\in A$ we have that
@@
We now analyze the left hand side of equation (131). As $\operatorname{\tilde{\mathcal{D}}}$ is empty, we find that
$\tilde{\chi}_{a}=\omega(T_{a}\tilde{\mathcal{M}}\cdot\Phi_{p}(\hat{\zeta}))\prod_{j=1}^{p-1}\lambda(T^{*}\tilde{F}_{a}\cdot\hat{\zeta}^{j})^{-1}.$ (132)
@@
$\operatorname{Td}\tilde{F}_{a}=\exp(c_{a}/2)\prod_{z}\frac{z}{e^{z/2}-e^{-z/2}}.$ (134)
@@
$\prod_{j=1}^{p-1}\operatorname{ch}(\lambda(T^{*}\tilde{F}_{a}\cdot\hat{\zeta}^{j})^{-1})=\prod_{\alpha\in\mu_{p}^{*}}\prod_{z}\frac{1}{1-\alpha e^{-z}}=\frac{\exp\left((p-1)c_{l}/2\right)}{\prod_{\alpha\in\mu_{p}^{*}}\prod_{z}\left(e^{z/2}-\alpha e^{-z/2}\right)}$ (135)
@@
$Z^{p}-Z^{-p}=\prod_{\alpha\in\mu_{p}}(Z-\alpha Z^{-1}),$ (137)
@@
$\prod_{j=1}^{p-1}\operatorname{ch}(\lambda(T^{*}\tilde{F}_{a}\cdot\hat{\zeta}^{j})^{-1})\cdot\operatorname{Td}\tilde{F}_{a}=p^{-d_{l}}\operatorname{Td}\tilde{F}_{a}(pz).$ (138)
@@ 3 Quantum Topology
Proof of Corollary LABEL:Cor:TopInv.
@@
It is thus clear that $(D_{\Lambda},l_{0})$ is determined by the Seifert invariants of $(M_{f},K)$ as defined in the introduction. ∎
We now explain the refinement formula given in equation (LABEL:eq:algquantumtop). We have that $\mathcal{N}^{f}$ is a subvariety of $\mathcal{M}^{\mathbb{T}\times\langle f\rangle}$. The torus $\mathbb{T}$ act trivially on the restriction of the determinant line bundle $\mathcal{L}$ to $\mathcal{N}$, whereas $\mathbb{T}$ acts with non-trivial weight $t^{(2-2g)m}$ on $\mathcal{L}$ restricted to $\mathcal{F}_{m}$ for each $m\in M$. Therefore the sum appearing in (LABEL:eq:Main) corresponding to components of $\mathcal{N}^{f}$ is seen to give the constant term of $\chi_{\mathbb{T}}(\mathcal{M},\mathcal{L}^{k},f)$. We have that $T_{\mathcal{N}}\mathcal{M}^{-}\cong T\mathcal{N}$, and $\det(T\mathcal{N})\cong\mathcal{L}^{2}$. As the rank of $T_{\mathcal{N}}\mathcal{M}^{-}$ is equal to 3g33g-3, the constant is equal to $(-1)^{3g-3}$ times $\chi(\mathcal{N},\mathcal{L}^{k+2},f)$. Thus we obtain equation (LABEL:eq:algquantumtop).
3.1 Complex Quantum Chern-Simons, TQFT Invariants and The Automorphism Equivariant Hitchin Index
Classical Chern-Simons theory with complex gauge group was introduced in [28], and complex quantum Chern-Simons was studied in the physics literature in [37, 75]. In [75], the moduli space of flat complex connections on a two-manifold was quantized with respect to a family of real polarizations (different from the one considered in this article). This quantization was subsequently studied by the first author et al in the works [4, 11]. A rigorous mathematical construction of a full TQFT modelling complex quantum Chern-Simons theory have yet to be given, although it is widely believed that the quantum Teichmüller TQFT [10, 9] of the first author and Kashaev is a part of it corresponding to the real gauge group $\operatorname{PSL}(2,\mathbb{R})$.
As explained in detail in the introduction, the study of the automorphism equivaraint index (LABEL:eq:index) is motivated by TQFT axioms [17] and the prospect of a full three-dimensional TQFT that models quantum Chern-Simons theory with complex gauge group $\operatorname{SL}(2,\mathbbm{C})$. In particular, it is natural to conjecture that the index (LABEL:eq:index) is closely connected to the partition function on the mapping torus with the presence of a (labelled) knot traced out by the marked point. This belief is supported by Corallary LABEL:Cor:TopInv. As such, Corollary LABEL:Cor:TopInv partially extends the bridge between algebraic geometry and quantum topology that exists via the realization of the TQFT representation of the mapping class group of a two-manifold through quantization of the moduli space of flat connections on said two-manifold. See e.g. [2, 15, 20, 48, 74] and the references therein.